A dye laser emits a pulse at 448 nm that contains 6.90 × 1015 photons. What is the energy content (in mJ) of this pulse?
Answer #1

Energy of the pulse is given by E = Nhc/ where N is number of photons, h is Planck's constant, wavelength and c is velocity of light. Therefore E = (6.90*10^15* 6.625*10^-34*3*10^8)/448*10^-9. = 0.306*10^-2 = 3.06*10-3 J = 3.06 mJ